# Mixture of Multicenter Experts in Multimodal Generative AI for Advanced Radiotherapy Target Delineation

### v1

@@ Abstract
Oncosoft Inc., Seoul, South Korea
These authors contributed equally
@@ Introduction
In this study, we applied MoME to address the limitations of current AI models in prostate cancer target volume contouring, that exhibits significant variability across institutions. Furthermore, we extend our approach to MoME fine-tuning by incorporating in-house fine-tuning for hospitals with restricted data sharing to address clinical deployment challenges when adapting AI models to new data distributions. Our results demonstrate that a MoME-based model not only significantly outperforms traditional AI models in target volume contouring, but also aligns its overall distribution more closely with that of each institution. This improvement is especially pronounced in cases with diverse data characteristics, highlighting the potential of the MoME approach to advance the adoption of AI in radiation oncology while addressing the challenges posed by subtle differences in institutional treatment protocols and patient distributions across each site.
@@ Results
Figure 1: Schematic of multicenter collaboration on multimodal AI training using our proposed Mixture of Multicenter Experts (MoME) approach. (a) Characteristics of each center, which can influence their radiotherapy target delineation tendencies, emphasize the need for an adaptive AI training approach. (b) Radiotherapy target delineation strategies for prostate cancer patients vary significantly across different centers. (c) Compared to traditional unipath single-center training, our multicenter MoME training leverages both shared and center-specific routing paths. These paths activate relevant expert modules tailored to the unique characteristics of each center given few-shot dataset. Moreover, where data export is restricted, a closed center MoME fine-tuning is employed, allowing for in-house model adaptation.
@@ 0.1 Center-specific radiotherapy target delineation strategy
We collected prostate cancer datasets from five centers and analyzed their distinct characteristics and target delineation tendencies, as illustrated in Fig.1(a). Centers A (Yonsei Cancer Center; YCC) and B (Yongin Severance; YS), both in South Korea, exhibit similar treatment protocols and image acquisition settings. Although Center B has a different National Comprehensive Cancer Network (NCCN)[22] risk group distribution compared to Center A, their treatment strategies remain largely similar, resulting in relatively homogeneous datasets. Center C (Gangnam Severance; GS), also based in South Korea, differs more from Centers A and B in terms of image acquisition settings, NCCN risk group distribution, and treatment strategies, although the racial distribution of patients is consistent. In contrast, Centers D (Massachusetts General Hospital; MGH) and E (MAYO Clinic), located in the USA, show greater variability, with substantial differences in national healthcare systems, NCCN risk group distribution, and treatment strategies. This increasing divergence across centers introduces significant challenges for model generalization, particularly when both imaging environments and clinical strategies differ substantially. These findings emphasize the need for adaptive approaches, such as the Mixture of Multicenter Experts (MoME), which can robustly handle cross-center variability. A detailed summary of the data distribution is provided in Supplementary Table1 and further discussed in Method section.
0.2 Multicenter Generalizable Multimodal MoME training
@@
We first trained baseline methods under traditional single-center AI training setting (Table 1(a)). Models trained exclusively on data from Center A exhibited overfitting to the training distribution, leading to suboptimal performance on Centers B and C dataset which have not seen by the model during training under the single-center setting, yielding Dice scores of 0.670 and 0.520 for each Center B and C. However, incorporating both imaging and textual data, as described in our previous study [4], the multimodal approach improved performance on external datasets compared with vision-only models with Dice scores of 0.732 and 0.597 for each Center B and C.
@@
Then, we performed experiments under the newly introduced multicenter AI training setting (Table 1(b)), which utilizes few-shot datasets from Centers B and C in addition to the training data from Center A. Despite this, vision-only models showed only marginal improvement over the single-center training setting with Dice scores of 0.680 and 0.561 for each Center B and C. In contrast, when trained with the vanilla multimodal AI, planning target volume (PTV) delineation performance significantly improved, particularly for Center C by showing 0.667 in Dice metric, where acquisition settings differed from those at Centers A and B. Moreover, When trained with our proposed MoME multimodal AI, PTV delineation performance improved further by demonstrating stable performance above 0.7 by achieving Dice scores of 0.747, 0.750 and 0.716 for each Center A, B and C. For comparison, we re-trained a publicly available 3D text-driven multimodal segmentation framework, ConTEXTualNet [23]. ConTEXTualNet demonstrated comparable performance to vanilla multimodal AI training, but still showed suboptimal performance for center with few-shot dataset with Dice scores of 0.721 and 0.663 for each Center B and C. The performance gap and statistical significance between vision-only and vanilla multimodal training method to our proposed MoME training method for each center dataset are visualized in Fig. 2(a).
@@
Table 1: Planning Target Volume (PTV) delineation performance for prostate cancer patients with Mixture of Multicenter Experts (MoME).
@@
Figure 2: Multicenter AI training comparison on radiotherapy target volume delineation. (a) Quantitative comparison under the multicenter AI training setting. (b) In an intermediate-risk patient case, the institution typically does not perform prophylactic nodal irradiation, yet both vision-only and multimodal AI erroneously included nodes in the delineation. In contrast, the multimodal MoME correctly focused on the prostate, excluding the nodes. (c) In another intermediate-risk, N0 case where none of the models included elective nodal irradiation, the institution generally defines a larger PTV margin around the prostate. The MoME model delineated the target volume with a larger margin, consistent with institutional practice, while both vision-only and multimodal AI applied smaller margins. (d) GTV-to-PTV ratio (GPR) distribution for (e-f) each Center A, B, and C patients on different training strategies.
@@
To systematically compare and quantify the impact of institutional practices on target delineation beyond individual examples, we introduced the GTV-to-PTV ratio (GPR), defined as the ratio of gross tumor volume (GTV) to PTV. This metric was proposed because it effectively captures two key factors where inter-institutional differences are likely to arise in prostate cancer radiotherapy: whether prophylactic nodal irradiatio (PNI) is performed and the extent of margin applied to the gross tumor volume. When nodal irradiation is performed, the pelvic lymph nodes are included in the PTV, leading to a lower GPR (Fig. 2(d)). Furthermore, the margin applied around the gross tumor also influences the GPR, making it a comprehensive measure that reflects both these key aspects of radiotherapy practice. To compare each institution’s approach, we exclusively analyzed N0 patients, who have been pathologically diagnosed with no metastasis in the lymph nodes, and excluded N1 patients, whose treatment planning might overlap across institutions.
We firstly compared the GPR distributions of the ground truth labels, vision-only AI, vanilla multimodal AI, and MoME models on test data from centers A, B, and C (Fig. 2(e), (f), and (g), respectively). The results from the ground truth GPR distribution for each dataset highlighted two distinct peaks around GPR values of 20% in the violin plot, corresponding to cases with higher GPR values and lower GPR values, respectively. This provides a clear visualization of the differences in clinical practices between the centers. Specifically, for Center A and B, a greater concentration of data was found in the lower GPR range, indicating that PNI is routinely performed in a significant proportion of patients, even in intermediate-risk cases. In contrast, the ground truth GPR distribution for Center C displayed a higher distribution of GPR values, indicating a lower frequency of PNI, particularly for intermediate-risk patients. This illustrates a critical divergence in clinical practice patterns, where Center A and B frequently employ PNI across very high, high, and intermediate-risk groups. Meanwhile, Center C tends to exclude RNI, particularly for intermediate-risk patients, reflecting a more conservative approach in this cohort.
@@
When stratified by risk group, the vision-only model consistently exhibited a high concentration of low GPR values across all risk levels, further confirming its inability to capture important clinical nuances. While the vanilla multimodal AI model showed some improvement, it still failed to accurately reflect the ground truth distribution. In contrast, the MoME model produced the most accurate GPR distribution, closely aligning with the original ground truth values across all risk groups (Supplementary Fig. 2). This highlights the MoME model’s superior capacity not only to incorporate clinical context but also to adapt to institution-specific practices, resulting in more precise and context-aware target delineation.
@@ 0.3 Center-specific Inference Reflects Radiotherapy Planning Strategy
Figure 3: Center-specific inference with the proposed multimodal MoME model. (a) Radiotherapy planning strategy change with different center-specific inference in GTV-to-PTV ratio (GPR) distribution. Given each center dataset as input, the corresponding center inference mostly reflects the planning target strategy. (b) Visualization of average frequency of activated top-k experts for each center dataset with corresponding center-specific router. (c) Visualization of frequency of activated top-k experts for an individual case. (d) Qualitative comparison of center-specific inference. Given a CT scan from Center C dataset, activating Center A and B-specific router results in a slightly enlarged PTV margin with prophylactic nodal irradiation, while the correct Center C-specific router aligns closely with the ground truth.
@@ 0.4 Data Efficient Few-shot Fine-tuning on Closed Center Dataset
To evaluate our proposed MoME training method in closed center settings, where data export is restricted, we conducted few-shot fine-tuning in-house. We adapted the pre-trained model to the newly introduced center dataset, leveraging the computing resources available within the center. To compare the training efficiency of each vision-only, vanilla multimodal, and MoME model, we monitored the performance of each method in PTV contouring as the size of the few-shot fine-tuning dataset progressively increased. Table 2 summarizes the performance of these methods across diverse closed center datasets.
Table 2: Planning Target Volume (PTV) delineation performance on closed center dataset with different size of few-shot fine-tuning dataset.
@@
For Center E, in the 0-shot inference setting as shown in Table 2, both vision-only and multimodal models exhibited limited effectiveness, with Dice scores of 0.327 and 0.476, respectively. In contrast, the MoME approach, when utilizing Center C-specific routers, achieved a higher Dice score of 0.558. This is notable because Center C shares the most similar data acquisition conditions with Center D, as detailed in Supplementary Table 1. Then, we further fine-tuned the network using Center E dataset. In the few-shot fine-tuning settings, the multimodal MoME consistently outperformed baseline models. Specifically, the MoME approach achieved Dice scores exceeding 0.6 across all datasets, with Dice scores of 0.630, 0.631, and 0.648 for the 1-shot, 2-shot, and 3-shot settings, respectively. The performance gains compared to vision-only and vanilla multimodal training mechanisms are also illustrated in Fig. 4(b). For both closed center fine-tuning experiments, our proposed MoME fine-tuning method demonstrated the generalizability and adaptability for distinct radiotherapy delineation strategy.
@@
Figure 4: Closed center dataset evaluation. (a) Closed Center D evaluation in Dice metric based on varying few-shot fine-tuning dataset sizes. (b) GTV-to-PTV ratio (GPR) distribution for each method, and MoME fine-tuning result with varying number of few-shot fine-tuning of the closed center D dataset. (c) Closed Center E evaluation in Dice metric based on varying few-shot fine-tuning dataset sizes. (d) GPR distribution comparison on closed center E dataset.
0.5 Ablation Study of Multicenter AI Training Strategy
@@
We conducted further ablation studies by replacing each module within our proposed MoME training strategy to assess the contribution of each component. First, to evaluate the multicenter training method, we designed different strategies to handle diverse data distribution: Text Prompt and Vanilla MoE methods. The Text Prompt method incorporated the center title, such as “Center C”, appended to the input clinical data within the baseline multimodal AI training framework. The Vanilla MoE method used a unified router for all center data, without a center-specific router. The results in Table. 3(a) compares these different training methods across three datasets. For both Center A and Center B, the Dice score and IoU were relatively consistent across the three methods, indicating no significant differences among them when applied to the primary training dataset or datasets with similar settings and distributions. In contrast, for Center C, our proposed MoME approach showed a significant improvement in Dice metric by achieving 0.716 compared to the Text Prompt and Vanilla MoE methods with Dice scores of 0.623 and 0.619, respectively. Moreover, MoME training achieved the lowest HD-95 value of 4.132 compared to 5.760 and 5.630 for each Text Prompt and Vanilla MoE, indicating its superior performance with reduced significant contouring errors. These results suggest that incorporating the center-specific router within our proposed MoME structure enhances adaptability during multi-center training, especially when substantial differences in data distribution exist among centers.
@@
We also evaluated the impact of varying the number of experts within the proposed MoME framework (ranging from Top-1 to Top-3), as shown in Table. 3(b). Our network consists of a total of 8 expert modules, with the Top-2 method being the baseline. Reducing the number of selected experts to Top-1 led to sparser center-specific training, while increasing it to Top-3 allowed greater overlap of experts across centers. The results showed that using Top-1 experts achieved the best performance for Centers A and B across all metrics, but its performance on Center C was significantly lower with Dice score of 0.674 than that of the Top-2 strategy with Dice score of 0.716. Meanwhile, Top-3 experts also demonstrated comparable performance for Center A and B, but the performance dropped for Centers C under 0.7 in Dice score. These results suggest that optimal performance in the allocation of experts requires balancing the number of experts in relation to the number of centers to avoid both excessive overlap and excessive sparsity in expert assignments, ensuring the MoME module is effectively adapted to each center-specific strategy.
Table 3: Ablation studies on network training strategy.
@@
Discussion
Since 2016, following the first FDA approval of a deep learning-based computer-aided diagnostic tool for clinical use [25], there has been a rapid expansion of research into artificial intelligence (AI) in medicine, with numerous AI models now integrated into clinical workflows [26]. While these advances have been promising, they are predominantly confined to applications like anomaly detection [27, 28] and organ segmentation [29, 30], areas where outputs are visually interpretable and unambiguous. However, the application of AI to more complex clinical tasks, such as treatment recommendations or interdisciplinary decision-making that require comprehensive multimodal inputs, remains underdeveloped. A major barrier to wider adoption is the significant variability in clinical practice guidelines across institutions and countries, which complicates the standardization needed for reliable AI-driven decision-making.
Two key challenges contribute to this limitation. The first is data dominance: even when AI models are trained on multi-institutional datasets, they often become biased toward the most prevalent patterns from dominant institutions. Consequently, the AI tends to recommend treatment approaches that reflect those institutions’ practices, which may not be appropriate or accepted in other medical settings. The second challenge is the considerable variation in clinical practices between institutions and countries, making it difficult for AI to propose universally standardized treatment plans. While it is possible to fine-tune AI models using data from individual institutions, the typical fine-tuning process requires a substantial amount of diverse data, which poses a significant barrier to quick adaptation across different institutions. This challenge requires the need for more flexible and adaptive models that can generalize better across varied clinical environments without the need for extensive retraining.
@@ Discussion
Radiation oncology is a field where these challenges are particularly pronounced. Despite the existence of international standards such as the National Comprehensive Cancer Network (NCCN) guidelines [22], expert panel consensus, and textbooks based on extensive clinical evidence, significant variability persists in how treatment protocols are applied. Differences in radiotherapy dose regimens, target structures, and institutional practices are notable. These variations are influenced by factors such as institutional resources, training methodologies, and even cultural factors, which complicate the establishment of a fully standardized approach across regions and clinical settings.

### v2

@@ Abstract
These authors contributed equally: Yujin Oh, Sangjoon Park, and Xiang Li
@@ Introduction
In this study, we applied MoME to address the limitations of current AI models in prostate cancer target volume delineation, that exhibits significant variability across institutions. Furthermore, we extend our approach to MoME fine-tuning by incorporating in-house fine-tuning for hospitals with restricted data sharing to address clinical deployment challenges when adapting AI models to new data distributions. Our results demonstrate that a MoME-based model not only significantly outperforms traditional AI models in target volume contouring, but also aligns its overall distribution more closely with that of each institution. This improvement is especially pronounced in cases with diverse data characteristics, highlighting the potential of the MoME approach to advance the adoption of AI in radiation oncology while addressing the challenges posed by subtle differences in institutional treatment protocols and patient distributions across each site.
@@ Results
Figure 1: Schematic of multimodal-multicenter AI training using our proposed Mixture of Multicenter Experts (MoME) approach. (a) Characteristics of each center, which can influence their radiotherapy target delineation tendencies, emphasize the need for an adaptive AI training approach. (b) Radiotherapy target delineation strategies for prostate cancer patients vary slightly across various centers, which limits the generalizability of AI models. (c) Compared to traditional unipath single-center training, our MoME training leverages both shared and center-specific routing paths. These paths activate relevant expert modules customized to the unique characteristics of each center given a few-shot dataset. Moreover, where data export is restricted, a closed center MoME fine-tuning is employed, allowing for in-house model adaptation.
@@ 0.1 Center-specific radiotherapy target delineation strategy
We collected prostate cancer datasets from five centers and analyzed their distinct characteristics and target delineation tendencies, as illustrated in Fig. 1(a). Centers A (Yonsei Cancer Center; YCC) and B (Yongin Severance; YS), both in South Korea, exhibit similar treatment protocols and image acquisition settings. Although Center B has a different National Comprehensive Cancer Network (NCCN) [22] risk group distribution compared to Center A, their treatment strategies remain largely similar, resulting in relatively homogeneous datasets. Center C (Gangnam Severance; GS), also based in South Korea, differs more from Centers A and B in terms of image acquisition settings, NCCN risk group distribution, and treatment strategies, although the racial distribution of patients is consistent. In contrast, Centers D (Massachusetts General Hospital; MGH) and E (MAYO Clinic) in the US show greater variability, with substantial differences in national healthcare systems, NCCN risk group distribution, and treatment strategies. This increasing divergence across centers introduces significant challenges for model generalization, particularly when both imaging environments and clinical strategies differ substantially. These findings emphasize the need for adaptive approaches, such as the Mixture of Multicenter Experts (MoME), which can robustly handle cross-center variability. A detailed summary of the data distribution is provided in Supplementary Table 1 and further discussed in the Methods section.
0.2 Analysis of model performance across centers
@@
We first trained baseline methods under the traditional single-center AI training setting (Table 1(a)). Models trained exclusively on data from Center A exhibited overfitting to the training distribution, leading to suboptimal performance on Centers B and C datasets, which were not seen by the model during training under the single-center setting, yielding Dice scores of 0.670 and 0.520 for Centers B and C, respectively. However, incorporating both imaging and textual data, as described in our previous study [4], the multimodal approach improved performance on external datasets compared with vision-only models, with Dice scores of 0.732 and 0.597 for Centers B and C, respectively.
@@
Then, we performed experiments under the newly introduced multicenter AI training setting (Table 1(b)), which utilizes few-shot datasets from Centers B and C in addition to the training data from Center A. Despite this, vision-only models showed only marginal improvement over the single-center training setting, with Dice scores of 0.680 and 0.561 for Centers B and C, respectively. In contrast, when trained with the vanilla multimodal AI, planning target volume (PTV) delineation performance significantly improved, particularly for Center C, showing 0.667 in Dice metric, where acquisition settings differed from those at Centers A and B. Moreover, when trained with our proposed MoME multimodal AI, PTV delineation performance improved further, demonstrating stable performance above 0.7 by achieving Dice scores of 0.747, 0.750, and 0.716 for Centers A, B, and C, respectively. For comparison, we re-trained a publicly available 3D text-driven multimodal segmentation framework, ConTEXTualNet [23]. ConTEXTualNet demonstrated comparable performance to vanilla multimodal AI training but still showed suboptimal performance for centers with few-shot datasets, with Dice scores of 0.721 and 0.663 for Centers B and C, respectively. The performance gap and statistical significance between vision-only and vanilla multimodal training methods compared to our proposed MoME training method for each center dataset are visualized in Fig. 2(a).
@@
Table 1: PTV delineation performance for prostate cancer patients with MoME.
@@
Figure 2: Multicenter AI training comparison on radiotherapy target volume delineation. (a) Quantitative comparison under the multicenter AI training setting. The center line indicates the median, the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile), and the x mark indicates the mean. The p-values indicate the statistically significant superiority of the proposed multimodal MoME to baseline methods. (b) In an intermediate-risk patient case, the institution typically does not perform prophylactic nodal irradiation, yet both vision-only and multimodal AI erroneously included nodes in the delineation. In contrast, the multimodal MoME correctly focused on the prostate, excluding the nodes. (c) In another intermediate-risk, N0 case where none of the models included elective nodal irradiation, the institution generally defines a larger PTV margin around the prostate. The MoME model delineated the target volume with a larger margin, consistent with institutional practice, while both vision-only and multimodal AI applied smaller margins. (d) GTV-to-PTV ratio (GPR) distribution. (e-g) GPR for each Centers A, B, and C patients under different training strategies. The center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).
@@
To systematically compare and quantify the impact of institutional practices on target delineation beyond individual examples, we introduced the GTV-to-PTV ratio (GPR), defined as the ratio of gross tumor volume (GTV) to PTV. This metric was proposed because it effectively captures two key factors where inter-institutional differences are likely to arise in prostate cancer radiotherapy: whether prophylactic nodal irradiation (PNI) is performed and the extent of margin applied to the gross tumor volume. When nodal irradiation is performed, the pelvic lymph nodes are included in the PTV, leading to a lower GPR (Fig. 2(d)). Furthermore, the margin applied around the gross tumor also influences the GPR, making it a comprehensive measure that reflects both these key aspects of radiotherapy practice. To compare each institution’s approach, we exclusively analyzed N0 patients, who have been pathologically diagnosed with no metastasis in the lymph nodes, and excluded N1 patients, whose treatment planning might overlap across institutions.
We first compared the GPR distributions of the ground truth labels, vision-only AI, vanilla multimodal AI, and MoME models on test data from centers A, B, and C (Figs. 2(e), (f), and (g), respectively). The results from the ground truth GPR distribution for each dataset highlighted two distinct peaks around GPR values of 20%, corresponding to cases with higher and lower GPR values, respectively. This provides a clear visualization of the differences in clinical practices between the centers. Specifically, for Centers A and B, a greater concentration of data was found in the lower GPR range, indicating that PNI is routinely performed in a significant proportion of patients, even in intermediate-risk cases. In contrast, the ground truth GPR distribution for Center C displayed a higher distribution of GPR values, indicating a lower frequency of PNI, particularly for intermediate-risk patients. This illustrates a critical divergence in clinical practice patterns, where Centers A and B frequently employ PNI across very-high, high, and intermediate-risk groups. Meanwhile, Center C tends to exclude PNI, particularly for intermediate-risk patients, reflecting a more conservative approach in this cohort.
@@
When stratified by risk group, the vision-only model consistently exhibited a high concentration of low GPR values across all risk levels, further confirming its limitations in capturing crucial clinical subtleties. While the vanilla multimodal AI model showed some improvement, it still failed to accurately reflect the ground truth distribution. In contrast, the MoME model produced the most accurate GPR distribution, closely aligning with the original ground truth values across all risk groups (Supplementary Fig. 2). This highlights the MoME model’s superior capacity not only to incorporate clinical context but also to adapt to institution-specific practices, resulting in more precise and context-aware target delineation.
@@ 0.3 Center-specific inference reflects radiotherapy planning strategy
Figure 3: Center-specific inference with the proposed multimodal MoME model. (a) Radiotherapy planning strategy change with different center-specific inference in GTV-to-PTV ratio (GPR) distribution. Given each center dataset as input, the corresponding center inference mostly reflects the planning target strategy. (b) Visualization of average frequency of activated top-k experts for each center dataset with corresponding center-specific router. (c) Visualization of frequency of activated top-k experts for an individual case. For (a-c), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile). (d) Qualitative comparison of center-specific inference. Given a CT scan from Center C dataset, activating Center A and B-specific router results in a slightly enlarged PTV margin with prophylactic nodal irradiation, while the correct Center C-specific router aligns closely with the ground truth.
@@ 0.4 Data efficient few-shot fine-tuning on closed center dataset
To evaluate the proposed MoME training method in closed center settings, where data export is restricted, we conducted few-shot fine-tuning in-house. We adapted the pre-trained model to the newly introduced center dataset, leveraging the computing resources available within the center. To compare the training efficiency of each vision-only, vanilla multimodal, and MoME model, we monitored the performance of each method in PTV contouring as the size of the few-shot fine-tuning dataset progressively increased. Table 2 summarizes the performance of these methods across diverse closed center datasets.
Table 2: PTV delineation performance on closed center dataset with different size of few-shot fine-tuning dataset.
@@
For Center E, in the 0-shot inference setting as shown in Table 2, both vision-only and multimodal models exhibited limited effectiveness, with Dice scores of 0.327 and 0.476, respectively. In contrast, the MoME approach, when utilizing Center C-specific routers, achieved a higher Dice score of 0.558. This is notable because Center C shares the most similar data acquisition conditions with Center E, as detailed in Supplementary Table 1. Then, we further fine-tuned the network using Center E dataset. In the few-shot fine-tuning settings, the multimodal MoME consistently outperformed baseline models. Specifically, the MoME approach achieved Dice scores exceeding 0.6 across all datasets, with Dice scores of 0.630, 0.631, and 0.648 for the 1-shot, 2-shot, and 3-shot settings, respectively. The performance gains compared to vision-only and vanilla multimodal training mechanisms are also illustrated in Fig. 4(b). For both closed center fine-tuning experiments, our proposed MoME fine-tuning method demonstrated generalizability and adaptability for distinct radiotherapy delineation strategies.
@@
Figure 4: Closed center dataset evaluation. (a) Closed Center D evaluation in Dice metric based on varying few-shot fine-tuning dataset sizes. (b) GTV-to-PTV ratio (GPR) distribution for each method, and MoME fine-tuning result with varying number of few-shot fine-tuning of the closed center D dataset. For (a-b), the Dice metric for each trial is presented as mean values (center lines) with 95th percentile of confidence intervals calculated with the non-parametric bootstrap method (shaded areas). (c) Closed Center E evaluation in Dice metric based on varying few-shot fine-tuning dataset sizes. (d) GPR distribution comparison on closed center E dataset. For (c-d), the center line indicates the median, and the bounds of the box represent the interquartile range (from the lower quartile to the upper quartile).
0.5 Ablation studies in MoME training strategy
@@
We conducted further ablation studies by replacing each module within the proposed MoME training strategy to assess the contribution of each component. First, to evaluate the multicenter training method, we designed different strategies to handle diverse data distribution: Text Prompt and Vanilla MoE methods. The Text Prompt method incorporated the center title, such as Center C, appended to the input clinical data within the baseline multimodal AI training framework. The Vanilla MoE method used a unified router for all center data, without a center-specific router. The results in Table 3(a) compare these different training methods across three datasets. For both Center A and Center B, the Dice score and IoU were relatively consistent across the three methods, indicating no significant differences among them when applied to the primary training dataset or datasets with similar settings and distributions. In contrast, for Center C, our proposed MoME approach showed a significant improvement in Dice metric by achieving 0.716 compared to the Text Prompt and Vanilla MoE methods with Dice scores of 0.623 and 0.619, respectively. Moreover, MoME training achieved the lowest HD-95 value of 4.132 compared to 5.760 and 5.630 for Text Prompt and Vanilla MoE, respectively, indicating its superior performance with reduced significant contouring errors. These results suggest that incorporating the center-specific router within our proposed MoME structure enhances adaptability during multi-center training, especially when substantial differences in data distribution exist among centers.
@@
We also evaluated the impact of varying the number of experts within the proposed MoME framework (ranging from Top-1 to Top-3), as shown in Table 3(b). Our network consists of a total of 8 expert modules, with the Top-2 method being the baseline. Reducing the number of selected experts to Top-1 led to sparser center-specific training, while increasing it to Top-3 allowed greater overlap of experts across centers. The results showed that using Top-1 experts achieved the best performance for Centers A and B across all metrics, but its performance on Center C was significantly lower with a Dice score of 0.674 compared to the Top-2 strategy with a Dice score of 0.716. Meanwhile, Top-3 experts also demonstrated comparable performance for Centers A and B, but the performance dropped for Center C to below 0.7 in Dice score. These results suggest that optimal performance in the allocation of experts requires balancing the number of experts in relation to the number of centers to avoid both excessive overlap and excessive sparsity in expert assignments, ensuring the MoME module is effectively adapted to each center-specific strategy.
Table 3: Ablation studies on the network training strategy.
@@
Discussion and Conclusion
Since 2016, following the first FDA approval of a deep learning-based computer-aided diagnostic tool for clinical use [25], research on AI applications in healthcare has grown rapidly, leading to the integration of various AI models into clinical settings [26]. Despite these advances, AI in clinical practice still faces persistent challenges, particularly in biases that arise from limited or skewed training data distributions [27, 28]. Specifically, many healthcare datasets underrepresent minority populations and lack critical demographic data, such as ethnicity and race [29]. For instance, a systematic review of cancer image datasets in dermatology revealed that limited population representation hinders these models’ applicability to diverse real-world clinical settings [28]. Additionally, in radiology, where data labeling by experts is standard, label inconsistency often arises due to subjective biases of individual clinicians, impacting the reliability of AI models [30]. Data distribution shifts further complicate model performance in real-world applications over time, as they can lead to reduced accuracy. Our Mixture of Multicenter Expert (MoME) framework is designed to tackle these challenges by creating tailored center-specific paths that utilize small, diverse samples to address inter-institutional variability. This approach aims to enhance model robustness across varied patient demographics and institutional approaches, potentially aligning AI tools more closely with real-world clinical diversity and evolving treatment philosophies.
@@
Radiation oncology is a field where these challenges are particularly pronounced. Despite the existence of international standards such as the National Comprehensive Cancer Network (NCCN) guidelines [22], expert panel consensus, and textbooks based on extensive clinical evidence, inter-clinician variability persists in how treatment protocols are applied. Differences in radiotherapy dose regimens, target structures, and institutional practices are notable. These variations are influenced by factors such as institutional resources, training methodologies, and even cultural factors, which complicate the establishment of a fully standardized approach across regions and clinical settings. To address the challenges of variability in patient characteristics and treatment strategies across different institutions and countries, we proposed the multicenter training strategy that incorporates diverse clinical approaches using MoME modules within a multimodal alignment learning framework. Our center-specific adaptation aims to manage the complexity and heterogeneity of patient populations and treatment protocols. This method allows the model to learn essential knowledge from large-scale datasets for patient treatment planning while simultaneously specializing in each center’s unique patient patterns with just one-shot training across five classes, requiring only 10 additional training samples per center, including a one-shot validation dataset.